A lady cannot see objects closer than 40 cm from the left eye and closer than 100 cm from the right eye. While on a mountaineering trip, she is lost from her team. She tries to make an astronomical telescope from her reading glasses to look for her teammates. (a) Which glass should she use as the eyepiece? (b) What magnification can she get with relaxed eye?

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. The lady can not see objects closer than 40 cm from the left eye and 100 cm from the right eye. For the left glass lens, v = – 40 cm, u = – 25 cm ∵ 1/f = 1/v – 1/u = 1/-40 – 1/-25 = 1/25 – 1/40 = 3/200 ⇒ f = 200/3 cm For the right glass lens, V = - 100 cm, u = - 25 cm 1/f = 1/v – 1/u = 1/-100- 1/-25 = 1/25 – 1/100 = 3/100 ⇒ f = 100/3 cm (a) For an astronomical telescope, the eye piece lens should have smaller focal length. So, she should use the right lens (f = 100/3 cm) as the eye piece lens. (b) With relaxed eye, (normal adjustment) F base 0 = 200/3 cm, f base e = 100/3 cm Magnification = m = f base 0/f base e = (200/3)/(100/3) = 2